# E2ENet: Dynamic Sparse Feature Fusion for Accurate and Efficient 3D Medical Image Segmentation

### v1

@@ Abstract
Deep neural networks have evolved as the leading approach in 3D medical image segmentation due to their outstanding performance. However, the ever-increasing model size and computation cost of deep neural networks have become the primary barrier to deploying them on real-world resource-limited hardware. In pursuit of improving performance and efficiency, we propose a 3D medical image segmentation model, named Efficient to Efficient Network (E2ENet), incorporating two parametrically and computationally efficient designs. i. Dynamic sparse feature fusion (DSFF) mechanism: it adaptively learns to fuse informative multi-scale features while reducing redundancy. ii. Restricted depth-shift in 3D convolution: it leverages the 3D spatial information while keeping the model and computational complexity as 2D-based methods. We conduct extensive experiments on BTCV, AMOS-CT and Brain Tumor Segmentation Challenge, demonstrating that E2ENet consistently achieves a superior trade-off between accuracy and efficiency than prior arts across various resource constraints. E2ENet achieves comparable accuracy on the large-scale challenge AMOS-CT, while saving over $68\%$ parameter count and $29\%$ FLOPs in the inference phase, compared with the previous best-performing method. Our code has been made available at: https://github.com/boqian333/E2ENet-Medical.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: A comparison of feature fusion schemes. The purple nodes depict features extracted from the backbone, while the green nodes depict the fused features. In particular, in DiNTS (d), red lines indicate information flow paths determined through neural architecture search techniques. In E2ENet (e), the red lines with different widths represent sparse information flows determined by the DSFF mechanism, allowing for efficient feature fusion. E2ENet is capable of dynamically learning how much of the features to fuse is derived from the backbone.
@@
3D medical image segmentation, plays an essential role in numerous clinical applications, including computer-aided diagnosis (Yu et al., 2020b) and image-guided surgery systems (Ronneberger et al., 2015), etc. Over the past decade, the rapid development of deep neural networks has achieved tremendous breakthroughs and significantly boosted the progress in this area (Zhou et al., 2018; Huang et al., 2020; Isensee et al., 2021). However, the model sizes and computational costs are also exploding, which deter their deployment in many real-world applications, especially for 3D models where resource consumption scales cubically (Hu et al., 2021; Valanarasu & Patel, 2022). This naturally raises a research question: Can we design a 3D medical image segmentation method that trades off accuracy and efficiency better subjected to different resource availability?
@@
Accurately segmenting organs is a challenging task in 3D medical image segmentation due to the variability in size and shape even among the same type of organ, caused by factors such as patient anatomy and disease stage. This diversity amplifies the difficulty to accurately identify and distinguish the boundaries of different organs, leading to potential errors in segmentation. One of the main solutions for accurate medical image segmentation is to favorably leverage the multi-scale features extracted by the backbone network, but this remains a long-standing problem. Pioneering work UNet (Ronneberger et al., 2015) utilizes skip connection to propagate detailed information to high-level features. More recently, UNet++ (Zhou et al., 2018), CoTr (Xie et al., 2021) and DiNTS (He et al., 2021) have developed more complex neural networks architectures (e.g. densely skip connections and attention mechanism) or optimization techniques (e.g. neural architecture search (NAS) (Elsken et al., 2019)) for cross-scale feature fusion. NAS initially searches for the network topology and operator (e.g., 3x3 Convolution and Maxpool) and subsequently optimize the model weights. However, NAS typically demands a significant amount of computational resources for exploring the network topology and assessing numerous candidate architectures, making it computationally expensive and time-consuming. For instance, C2FNAS (Yu et al., 2020a) necessitates nearly one GPU year to discover a single 3D segmentation architecture, DiNTS (He et al., 2021) improves the searching efficiency but still requires 5.8 GPU days to search for a single architecture. Compared to NAS approaches, our method does not require costly architecture search time and directly optimizes/searches sparse topology within the pre-defined architecture.
@@
In this paper, we propose Efficient to Efficient Network (E2ENet), a model that can efficiently incorporate both bottom-up and top-down information flows from the backbone network in a dynamic sparse pattern, achieving a much better accuracy-efficiency trade-off. As shown in Figure 1 (e), E2ENet incorporates multi-scale features from the backbone into the final output by gradually fusing the adjacent features, allowing the network to fully utilize information from various scales. To prevent unnecessary information aggregation, a dynamic sparse feature fusion (DSFF) mechanism is proposed and embedded in each fusion node. The DSFF mechanism adaptively integrates relevant multi-scale features and filters out unnecessary ones during the course of the training process, significantly reducing the computational overhead without sacrificing performance. Additionally, to further improve efficiency, our E2ENet employs a restricted depth-shift strategy in 3D convolutions, which is derived from temporal shift (Lin et al., 2019) and 3D-shift (Fan et al., 2020) used in efficient video action recognition. This allows the 3D convolution operation with kernel size $(1,3,3)$ to capture 3D spatial relationships while maintaining the computation and parameter complexity of 2D convolutions.
To evaluate the performance of our proposed E2ENet, we conducted extensive experiments on the three medical image segmentation challenges, including BTCV, AMOS-CT and Brain Tumor Segmentation Challenge. E2ENet can effectively trade-off between segmentation accuracy and model efficiency compared with both convolution-based and transformer-based architectures. In particular, as shown in Figure 2, on the AMOS-CT challenge, E2ENet achieves competitive accuracy with mDice of 90.1%, while being 69% smaller and using 29% fewer FLOPs in the inference phase. Furthermore, experiments demonstrate that even when filtering out 90% of connections during feature fusion by the DSFF mechanism, E2ENet can maintain a comparable mDice performance with only a minor impact.
Figure 2: Performance comparison of E2ENet with three different levels of sparsity $0.9$, $0.8$ and $0.7$ and other models on the AMOS-CT challenge in terms of mDice (%), Params (M), and inference FLOPs (G). The feature sparsity level, denoted by $S$, is explained in Section 3.2.
2 Related Work
2.1 3D Medical Image Segmentation
Convolutional neural networks (CNNs) have become the dominant architecture for 3D medical image segmentation in recent years (e.g. 3D UNet (Çiçek et al., 2016), UNet++ (Zhou et al., 2018), UNet3+ (Huang et al., 2020), PaNN (Zhou et al., 2019) and nnUNet (Isensee et al., 2021)), due to their ability to capture local and weight-sharing dependencies (d’Ascoli et al., 2021; Dai et al., 2021). However, some recent methods have attempted to incorporate transformer modules into CNNs (e.g. CoTr (Xie et al., 2021), TransBTS (Wang et al., 2021)), or use pure transformer architectures (e.g. ConvIt (Karimi et al., 2021), nnFormer (Zhou et al., 2021), Swin UNet (Cao et al., 2021)), in order to capture long-range dependencies. These transformer-based approaches often require large amounts of training data, longer training times, or specialized training techniques, and can also be computationally expensive. In this paper, we propose an alternative method for efficiently incorporating 3D contextual information using a restricted depth-shift strategy in 3D convolutions, and further improving performance through adaptive multi-scale feature fusion.
2.2 Feature Fusion in Medical Image Segmentation
Multi-scale feature fusion is a crucial technique in medical image segmentation that allows a model to detect objects across a range of scales, while also recovering spatial information that is lost during pooling (Wang et al., 2022; Xie et al., 2021). However, effectively representing and processing multi-scale hierarchy features can be challenging, and simply summing them up without distinction can lead to semantic gaps and degraded performance (Wang et al., 2022; Tan et al., 2020). To address this issue, various approaches have been proposed, including adding learnable operations to reduce the gap with residuals (Ibtehaz & Rahman, 2020), attention blocks (Oktay et al., 2018). More recently, UNet++ (Zhou et al., 2018) and its variants (Li et al., 2020; Huang et al., 2020; Jha et al., 2019) have adapted the gating signal to dense nesting levels, taking into account as many feature levels as possible. NAS-UNet (Weng et al., 2019) tries to automatically search for better feature fusion topology. While these methods have achieved better performance, they can also incur significant computational and information redundancy. Dynamic convolution (Su et al., 2020; Chen et al., 2020) utilizes coefficient prediction or attention modules to dynamically aggregate convolution kernels, thereby reducing computation costs. In our paper, we propose an intuitive approach to optimizing multi-scale feature fusion, which enables selective leveraging of sparse feature representations from fine-grained to semantic levels through the proposed dynamic sparse feature fusion mechanism.
2.3 Sparse Training
Recently, sparse training techniques have shown the possibility of training an efficient network with sparse connections that match (or even outperform) the performance of dense counterparts with lower computational cost (Mocanu et al., 2018; Liu et al., 2021b). Beginning with (Mocanu et al., 2016), it has been demonstrated that initializing a static sparse network without optimizing its topology during training can also yield comparable performance in certain situations (Lee et al., 2018; Tanaka et al., 2020; Wang et al., 2019). However, Dynamic Sparse Training (DST), also known as sparse training with dynamic sparsity (Mocanu et al., 2018), offers a different approach by jointly optimizing the sparse topology and weights during the training process starting from a sparse network (Liu et al., 2021a, 2022; Evci et al., 2020; Jayakumar et al., 2020; Mostafa & Wang, 2019; Yuan et al., 2021). This allows the model’s sparse connections to gradually evolve in a prune-and-grow scheme, leading to improved performance compared to naively training a static sparse network (Liu et al., 2021c; Xiao et al., 2022). In contrast to prior methods that aim to find sparse networks that can match the performance of corresponding dense networks, we aim to leverage DST to adaptively fuse multi-scale features in a computationally efficient manner for 3D medical image segmentation.
@@ 3 Methodology
Figure 3: The overall architecture of the proposed E2ENet model consists of a CNN backbone that extracts multiple levels of features. These features are then gradually aggregated through several stages, during which the multi-scale features are fused using a fusion operation. The model has three main flow pathways: Downward flow (yellow) provides details from a lower to a higher level by down-sampling the features; Forward flow (red) applies multi-scale features fusion while maintaining resolution; Upward flow (blue) provides semantic information from a higher level to lower levels.
@@ 3.1 The Architecture
Figure 3 provides an overview of the proposed architecture. Given the input 3D image $I_{in}\in\mathbb{R}^{D\times H\times W}$, the CNN backbone extracts feature maps at multiple scales, represented by ${\mathbf{x}}^{0}=\left(\mathbf{x}^{0,1},\mathbf{x}^{0,2},\ldots,\mathbf{x}^{0,L}\right)$, where $L$ is the total number of feature scales. The feature at level ${i}$, $\mathbf{x}^{0,i}$ is a tensor with dimensions $d_{i}\times h_{i}\times w_{i}\times c_{i}$, where $d_{i},h_{i},w_{i}$ and $c_{i}$ represent the depth, height, width, and number of channels of the feature maps at level $i$, respectively. It’s worth mentioning that the spatial resolution of the feature maps decreases as the level increases, while the number of channels increases to a maximum of $320$. The backbone generates a total of $L=6$ multi-scale feature levels, each with a specified number of channels: [c1, c2, c3, c4, c5, c6] = [48, 96, 192, 320, 320, 320]. At each level of feature generation, there are two convolution layers with a kernel size of (1, 3, 3), followed by instance normalization and the application of leaky ReLU activation. The down-sampling ratios for each level are as follows: ((1, 2, 2), (2, 2, 2), (2, 2, 2), (2, 2, 2), (2, 2, 2), (2, 2, 2), (2, 2, 2)).
To fully exploit the hierarchical features extracted by the CNN backbone, they are aggregated through multiple stages, as depicted in Figure 3 (green part). At stage $j$ ($10<j\leq L-$ 1) of this process, the features at level $i$ in the current stage are the result of fusing the adjacent features from the previous stage along three directions: 1. “Downward flow” (in yellow): The high-resolution feature $\mathbf{x}^{j-1,i-1}$, which provides richer visual details, is passed downward; 2. “Upward flow” (in blue): The low-resolution feature $\mathbf{x}^{j-1,i+1}$, which captures more global context, is passed upward; 3. “Forward flow” (in red): The features $x^{j-1,i}$, which maintain their spatial resolution, are passed forward for further information integration. The fused cross-scale feature map at the $i$-th level of the $j$-th stage can be formulated as:
@@
where $\mathcal{F}^{j,i}(.)$ is a fusion operation, consisting of a 3D convolution operation followed by Instance Normalization (IN) (Ulyanov et al., 2016) and a Leaky Rectified Linear Unit (LeakyReLU) (Maas et al., 2013). $\mathcal{U}(.)$ and $\mathcal{D}(.)$ denote the up-sampling (transposed convolution with stride 2) and down-sampling (maxpooling), respectively. [.][.][ . ] denotes the concatenation operation. In the 3D convolution operation, the input feature maps, which have a channel number of $C^{j,i}_{in}$ $(C^{j,i}_{in}=c_{i-1}+c_{i}+c_{i+1}$ or $c_{i}+c_{i+1})$, are fused and processed to produce output feature maps with a channel number of $C^{j,i}_{out}$ $(C^{j,i}_{out}=c_{i})$.
We have integrated a deep supervision strategy to improve the training process, as in (Isensee et al., 2021). This strategy involves using downsampled versions of the ground truth segmentation and the corresponding predictions at each scale $i$ to compute the loss $l_{i}$. As shown in Figure 3, multi-scaled feature maps are processed through a $1\times 1\times 1$ convolutional layer in the output module, to generate multi-scaled predictions. The overall training objective is the sum of the losses at all scales, $l=w_{1}\times l_{1}+w_{2}\times l_{2}+…$, where the loss weights $w_{i}$ are normalized to sum to $1$ and decrease geometrically with the resolution.
Unlike the UNet++ model, which only considers bottom-up information flows for image segmentation, our proposed E2ENet architecture incorporates both bottom-up and top-down information flows. This allows E2ENet to make use of both high-level contextual information and fine-grained details in order to produce more accurate segmentation maps (experimental results can be found in Table 3).
@@ 3.2 Dynamic Sparse Feature Fusion
Such multistaged cross-level feature propagation provides a more comprehensive understanding of the images, but can also introduce redundant information, which necessitates careful treatment in feature fusion to ensure efficient interaction between scales or stages. Our proposed DSFF mechanism addresses the issue of multi-scale feature fusion in an intuitive and effective way. It optimizes the process by allowing selective and adaptive use of features from different levels during training. This results in a more efficient feature fusion process with lower computational and memory overhead.
The DSFF mechanism is applied in each fusion operation, allowing the fusion operation $\mathcal{F}^{j,i}(\cdot)$ to select the informative feature maps from the input fused features. The feature map selection process is controlled by a binary mask $\mathbf{M}^{j,i}(\cdot)\in\{0,1\}^{C^{j,i}_{in}\times C^{j,i}_{out}}$, which are trained to filter out $S$ unnecessary feature map connections by zeroing out all the values of the corresponding kernels in $\mathcal{F}^{j,i}(\cdot)$. With DSFF mechanism, the output of the fusion operation is then computed as:

### v2

@@ Abstract
Deep neural networks have evolved as the leading approach in 3D medical image segmentation due to their outstanding performance. However, the ever-increasing model size and computational cost of deep neural networks have become the primary barriers to deploying them on real-world, resource-limited hardware. To achieve both segmentation accuracy and efficiency, we propose a 3D medical image segmentation model called Efficient to Efficient Network (E2ENet), which incorporates two parametrically and computationally efficient designs. i. Dynamic sparse feature fusion (DSFF) mechanism: it adaptively learns to fuse informative multi-scale features while reducing redundancy. ii. Restricted depth-shift in 3D convolution: it leverages the 3D spatial information while keeping the model and computational complexity as 2D-based methods. We conduct extensive experiments on AMOS, Brain Tumor Segmentation and BTCV Challenge, demonstrating that E2ENet consistently achieves a superior trade-off between accuracy and efficiency than prior arts across various resource constraints. E2ENet achieves comparable accuracy on the large-scale challenge AMOS-CT, while saving over $69\%$ parameter count and $27\%$ FLOPs in the inference phase, compared with the previous best-performing method. Our code has been made available at: https://github.com/boqian333/E2ENet-Medical.
@@ 1 Introduction
3D medical image segmentation plays an essential role in numerous clinical applications, including computer-aided diagnosis (Yu et al., 2020) and image-guided surgery systems (Ronneberger et al., 2015). Over the past decade, the rapid development of deep neural networks has achieved tremendous breakthroughs and significantly boosted the progress in this area (Zhou et al., 2018; Huang et al., 2020; Isensee et al., 2021). However, the model sizes and computational costs are also exploding, which deter their deployment in many real-world applications, especially for 3D models where resource consumption scales cubically (Hu et al., 2021; Valanarasu and Patel, 2022). This naturally raises a research question: Can we design a 3D medical image segmentation method that trades off accuracy and efficiency better, subjected to different resource availability?
Figure 1: A comparison of feature fusion schemes. The purple nodes depict features extracted from the backbone, while the green nodes depict the fused features. In particular, in DiNTS (d), red lines indicate information flow paths determined through neural architecture search techniques. In E2ENet (e), the red lines with different widths represent sparse information flows determined by the DSFF mechanism, allowing for efficient feature fusion. E2ENet is capable of dynamically learning how many of the features to fuse are derived from the backbone.
@@
Accurately segmenting organs is a challenging task in 3D medical image segmentation due to the variability in size and shape even among the same type of organ, caused by factors such as patient anatomy and disease stage. This diversity amplifies the difficulty to accurately identify and distinguish the boundaries of different organs, leading to potential errors in segmentation. One of the main solutions for accurate medical image segmentation is to favorably leverage the multi-scale features extracted by the backbone network, but this remains a long-standing problem. Pioneering work UNet (Ronneberger et al., 2015) utilizes skip connection to propagate detailed information to high-level features. More recently, UNet++ (Zhou et al., 2018), CoTr (Xie et al., 2021), and DiNTS (He et al., 2021) have introduced more complex neural network architectures (e.g., dense skip connection and attention mechanism) and optimization techniques (e.g., neural architecture search (NAS) (Elsken et al., 2019)) for cross-scale feature fusion. NAS-based methods search for network topology and operators (e.g., 3x3 Convolution and Maxpool) and subsequently optimize model weights. However, these methods typically require substantial computational resources to explore network topologies and evaluate numerous candidate architectures, making them time-consuming. For instance, C2FNAS (Yu et al., 2020) requires nearly one GPU-year to discover a single 3D segmentation architecture, while DiNTS (He et al., 2021) improves search efficiency but still needs 5.8 GPU days to find a single architecture. In contrast to NAS approaches, our method does not require costly architecture search time and instead directly optimizes and searches for sparse topologies within the predefined architecture.
@@
In this paper, we propose the Efficient to Efficient Network (E2ENet), a model that efficiently incorporates both bottom-up and top-down features from the backbone network in a dynamically sparse pattern, achieving an improved accuracy-efficiency trade-off. As shown in Figure 1 (e), E2ENet incorporates multi-scale features from the backbone into the final output by gradually fusing adjacent features, allowing the network to fully utilize information across various scales. To prevent unnecessary information aggregation, we propose a dynamic sparse feature fusion (DSFF) mechanism and embed it in each fusion node. The DSFF mechanism adaptively integrates informative multi-scale features and filters out unnecessary ones during the course of the training process, significantly reducing the computational overhead without sacrificing performance. Additionally, to further improve efficiency, our E2ENet employs a restricted depth-shift strategy within 3D convolutions, derived from the temporal shift (Lin et al., 2019) and 3D-shift (Fan et al., 2020) strategies used in efficient video action recognition. This allows the 3D convolution operation with a kernel size of $(1,3,3)$ to capture 3D spatial relationships while maintaining the parameter complexity of 2D convolutions and reducing computational cost. To evaluate the performance of our proposed E2ENet, we conducted extensive experiments on the AMOS-CT (Ji et al., 2022), Brain Tumor Segmentation in the Medical Segmentation Decathlon (MSD) (Antonelli et al., 2022), and Multi-Atlas Labeling Beyond the Cranial Vault (BTCV) (Landman et al., 2015) challenges. We found that E2ENet can effectively trade-off between segmentation accuracy and efficiency compared to both convolution-based and transformer-based architectures. In particular, on the AMOS-CT challenge, E2ENet achieves competitive accuracy with a mDice of 90.0%, while being 69% smaller and using 27% fewer FLOPs during inference. With the DSFF mechanism filtering out 90% of feature connections, E2ENet further reduces resource costs without significantly sacrificing accuracy.
@@ 2 Methodology
Figure 2: The overall architecture of the proposed E2ENet consists of a CNN backbone that extracts multiple levels of features. These features are then gradually aggregated through several stages, during which the multi-scale features are fused using a fusion operation.
@@ 2.1 The Architecture
Figure 2 provides an overview of the proposed architecture. Given the input 3D image $I_{in}\in\mathbb{R}^{D\times H\times W}$, the CNN backbone extracts feature maps at multiple scales, represented by ${\mathbf{x}}^{0}=\left(\mathbf{x}^{0,1},\mathbf{x}^{0,2},\ldots,\mathbf{x}^{0,L}\right)$, where $L$ is the total number of feature scales. The feature at level ${i}$, $\mathbf{x}^{0,i}$ is a tensor with dimensions $d_{i}\times h_{i}\times w_{i}\times c_{i}$, where $d_{i},h_{i},w_{i}$ and $c_{i}$ represent the depth, height, width, and number of channels of the feature maps at level $i$, respectively. It is worth mentioning that the spatial resolution of the feature maps decreases as the level increases, while the number of channels increases up to a maximum of $320$. To fully exploit the hierarchical features extracted by the CNN backbone, these features are aggregated across multiple stages, as depicted in Figure 2 (Dynamic Sparse Feature Fusion section). At stage $j$ ($10<j\leq L-$ 1), the features at level $i$ are obtained by fusing the adjacent features from the previous stage along three directions:
1. “Downward flow” (in yellow): The high-resolution feature $\mathbf{x}^{j-1,i-1}$, which provides richer visual details, is passed downward; 2. “Upward flow” (in blue): The low-resolution feature $\mathbf{x}^{j-1,i+1}$, which captures more global context, is passed upward; 3. “Forward flow” (in red): The features $x^{j-1,i}$, which maintain their spatial resolution, are passed forward for further information integration. The fused cross-scale feature maps at the $i$-th level of the $j$-th stage can be formulated as:
@@
where $\mathcal{F}^{j,i}(.)$ is a fusion operation, consisting of a convolution operation followed by Instance Normalization (IN) (Ulyanov et al., 2016) and a Leaky Rectified Linear Unit (LeakyReLU) (Maas et al., 2013). $\mathcal{U}(.)$ and $\mathcal{D}(.)$ denote the up-sampling and down-sampling, respectively. [.][.][ . ] denotes the concatenation operation. In the convolution operation, the input feature maps, which have a channel number of $C^{j,i}_{in}$ $(C^{j,i}_{in}=c_{i-1}+c_{i}+c_{i+1}$ or $c_{i}+c_{i+1})$, are fused and processed to produce output feature maps with a channel number of $C^{j,i}_{out}$ $(C^{j,i}_{out}=c_{i})$. Unlike the UNet++ model, which only considers bottom-up information flows for image segmentation, our proposed E2ENet architecture incorporates both bottom-up and top-down information flows. This allows E2ENet to make use of both high-level contextual information and fine-grained details in order to produce more accurate segmentation maps (experimental results can be found in Table 2).
@@ 2.2 Dynamic Sparse Feature Fusion
Such multi-stage cross-level feature propagation provides a more comprehensive understanding of the images but can also introduce redundant information, necessitating careful handling in feature fusion to ensure efficient interaction between scales or stages. Our proposed Dynamic Sparse Feature Fusion (DSFF) mechanism addresses the issue of multi-scale feature fusion in an intuitive and effective way. It optimizes the process by enabling selective and adaptive use of features from different levels during training. This results in a more efficient feature fusion process with lower computational and memory overhead.
The DSFF mechanism is applied in each fusion operation, allowing the fusion operation $\mathcal{F}^{j,i}(\cdot)$ to select the informative feature maps from the input fused features. The feature map selection process is controlled by a binary mask $\mathbf{M}^{j,i}(\cdot)\in\{0,1\}^{C^{j,i}_{in}\times C^{j,i}_{out}}$, which are trained to filter out $S$ unnecessary feature map connections by zeroing out the corresponding kernels in $\mathcal{F}^{j,i}(\cdot)$. With the DSFF mechanism, the output of the fusion operation is then computed as: